OpenStudy (anonymous):

Find dy/dx of y=x(lnx)^(sin^2x) using log differentiation. lny = ln x + (sin^2 x) ln (lnx) y`/y = 1/x + sin2x (ln(lnx)) + sin^2x/(x lnx) Can someone explain to be thoroughly the lny = ln x + (sin^2 x) ln (lnx) step please!!!!! Test tomorrow

OpenStudy (rational):

y=x(lnx)^(sin^2x) take ln both sides, what do you get ?

OpenStudy (anonymous):

lny = .... ln(x(lnx)^(sin^2x))

OpenStudy (anonymous):

But wouldn't that =

OpenStudy (anonymous):

lny = (sin^2x)ln(x)(lnx) ? I'm messing up badly on this step lol

OpenStudy (perl):

$$ \Large y=x~ln(x)^{sin^2x} \\ \text{take ln of both sides} \\ \Large ln(y) = ln ( x~ln(x)^{sin^2x}) $$

OpenStudy (anonymous):

My eyes tell me that this is the log(a*b) rule but I don't see that in the answer either

OpenStudy (perl):

$$ \Large y=x\cdot ln(x)^{sin^2x} \\ \text{take ln of both sides} \\ \Large ln(y) = ln ( x\cdot ln(x)^{sin^2x}) = ln ( x) + ln( ln(x)^{sin^2x}) $$

OpenStudy (anonymous):

Oh okay I see the mult rule now, cool beans

OpenStudy (anonymous):

And then the power is pulled down afterwards

OpenStudy (perl):

$$ \Large \color{blue} {\text{Log rules:}}\\ \Large \log(A \cdot B) = \log(A)+ ln(B)\\ \Large \log(\frac{A}{B}) = \log(A)- ln(B)\\ \Large \log(A^n) = n\cdot \log(A) $$

OpenStudy (anonymous):

Thank ya this helps me. I have a couple more questions to clear up for my test if you want to have some more calc fun look for my questions! :D

OpenStudy (perl):

sure :)